# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ 10.1.4 Further remarks
Although the post-filtering has been applied with great caution, it also happens that some problems persist. One of these cases is the following. Some objects exhibit a good SB1 or SB1C solution with a well established significant orbital semi-amplitude for Gaia. However, these solutions could sometimes be doubtful because the same object observed in the framework of the APOGEE DR17 survey (Abdurro’uf et al. 2022) or LAMOST survey (Cui et al. 2012; Bashi et al. 2022) appears to be constant in RV as detailed in Sect. 8.2.4. This anomaly has not been fully explained but it could be related in some cases to the presence of a nearby neighbour (see below). These objects have not been rejected. We encourage the catalogue user to be very cautious when using the present SB-subcatalogue despite the fact that persisting problems should only represent a minority. All the remaining and identified problems should be corrected in the future DR4.
@@ 10.2 An explanation for Gaia DR3 3376949338201650112
In Sect. 8.2.4, we report the existence of SB1 solutions presenting a marked periodic variation, whereas ground-based data conclude to the constancy of the related object. Here below, we propose an explanation for this discrepancy by investigating in details the case of Gaia DR3 3376949338201650112. This example has the advantage to be enlightening and will allow us to correct the pipeline for future releases.
@@
The astrometric solution for this object is characterised by a ruwe of 25.35, pinpointing to a problem with the fit. Referring to the IPD parameters (Image Parameter Determination, see e.g. Holl et al. 2023a), we remark that ipd_gof_harmonic_amplitude is equal to 0.164 which is certainly different from zero. This is a clue that the object is badly treated as a single object. The value of ipd_frac_multi_peak turns out to be 84 suggesting that the object is expected to be double or multiple on the plane of the sky: most probably a resolved or, more precisely, a partially resolved pair (according to the definition of Holl et al. 2023a). As explained by these authors, such pairs (or higher order) have a predisposition to suffer from scan-angle effects that induce a bias on the measured RVs that is function of the relative orientation of the scanning direction on the sky at the time of the observations. This is confirmed for Gaia DR3 3376949338201650112 by Fig. 50 where it becomes clear that the measured RVs are function of the scan angle. The parameter ipd_gof_harmonic_phase adopts a value of 103.4 ° that suggests that the line joining the two stars of the pair has a position angle on the sky of 13.4 °, after the application of a 90 ° correction as explained in section 3.1.2 of Holl et al. (2023a). The semi-amplitude of the fake orbital solution is about 16.6 km s-1 which translates according to equation 2 of Holl et al. (2023a) into a separation of the pair of 114 mas. The existence of such a pair is confirmed by a preliminary run of AGIS (AGIS 4.1) linked to the future DR4 that actually detects two objects Gaia DR4 3376949342501945344 and Gaia DR4 3376949342501945472 that are separated by some 200 mas and are exhibiting an orientation of 9 °, thereby supporting our interpretation.
@@
In principle, the bias due to the scan-angle effect is intended to produce variability with a period either of 62.97 d or of 1 y. This property has been used in Sect. 10.1.3 to apply post-filtering. However, in the present case, the period is 0.657582 d which does not appear to be compatible and could shed some doubt on our interpretation. The spectral window associated to the DR3 RV time-series corresponding to this object is given in Fig. 51. Besides the main peak at d-1 as usual, it exhibits an important peak at d-1 that is exceeding a value of 0.8; this is an indication that isolated aliasing could be important for this particular time-series. An aliasing with d-1 explains the ambiguity between a period of 0.657582 d and one of one year. The combination of the aliasing phenomenon and of the scan-angle effect explains the existence of such fake SB1 objects in our SB-subcatalogue: such objects were not post-filtered, since the post-filtering was based on the 62.97 d or 1 y timescales. We conclude that it would be more appropriate to include a statistical test inspired by graphs like the one shown in Fig. 50 rather than on a criterion based on the period. However, a true periodicity with these timescales will imply a similar behaviour of the RV dependency on the scan angle even if no scan-angle effect is present.
@@ 10.3 The case of astrometric binaries
In Sect. 10.1.3 and in Sect. 10.2, we discussed the effect of astrometric close pairs on the determination of RVs. A close pair fixed on the sky could often generate a scan-angle effect and lead to corrupted RV values but also could generate fake orbital motions. An additional degree of complexity is reached if one considers astrometric binaries. Due to their binarity, these close pairs are not fixed on the sky and the main axis of the pair is rotating due to the orbital motion. Therefore, there is interaction between the above-mentioned scan-angle effect and the astrometric orbital motion itself. The situation still increases in complexity if the astrometric orbital motion and the spectroscopic RV changes correspond to two aspects of the very same orbital motion and thus have the same period. In this latter case, the correction presents a timescale equal to the phenomenon we are studying. The correcting scheme is explained in section 3.3.1 of Holl et al. (2023a) and is also illustrated there with the (somewhat extreme) example of Gaia DR3 6631710607341412096. For details, we refer to this paper.
@@ 11.1 Description of the output
The present first catalogue of pure spectroscopic solutions is part of the Gaia Data Release 3. The orbital solutions (in class SB1 or SB1C as indicated in the entry nss_solution_id) are part of an auxiliary table named gaiadr3.nss_two_body_orbit. We output in this table the classical Campbell orbital parameters introduced in Sect. 3.1 for the SB1 class (, , , , , ) along with their estimated uncertainties (1). The same is done for circular orbit parameters (, , , ) introduced in Sect. 3.2 and corresponding to the SB1C class. Beyond the orbital parameters themselves, this table also contains some indicators of the quality of the solution (see Sect. 6.4). The flags associated to the solution (see Sect. 6.5) are also provided. The boolean mask bit_index defines the parameters that are actually fitted, and corresponds for the SB1 class (respectively SB1C) to the value 127 (respectively 31). In addition, is also provided the upper triangle of the correlation matrix of the parameters. It is expressed as corr_vec that contains the cross-coefficients folded into a vector and corresponding to the ordering of the adopted list of parameters. The first term of the vector gives the element (0,1) of the matrix, the element gives the element (0,2) of the matrix and element the (1,2) one. Elements , , correspond respectively to the matrix elements (0,3), (1,3) and (2,3) and so on. This is thus a column-major storage scheme. The details of the output are described in Table 15 for the class SB1 and in Table 16 for the SB1C one. The table contains 181327 objects (with solutions) in the SB1 class and 202 in the SB1C one. The marked difference in the counting is due to our decision to consider eccentric solutions down to as small eccentricities as possible before going circular. This was described in Sect. 6.3.
@@
11.2 Weaknesses, caveats and recommendations
@@
Another case of trickery is the one of true SB2 stars, where the two lines never fully deblend. They are appearing most probably in the SB1 channel. The RV variations are then due to the corresponding change of the full profile made by the blended lines. This has little effect on the period determination but the impact on the eccentricity and mainly on the semi-amplitude might be huge (Sect. 8.2.3). The related solutions are characterised by a too small value of and of the mass function. They could be considered as suspicious but we can not have any certitude that all of these types of solutions are necessarily wrong.
@@
Instrumental effects are also present. The scan-angle effect notably induces periodic RV variations for constant stars that do not fulfil the assumptions used in astrometry (objects are assumed to be isolated and thus single). Close astrometric pairs could thus mimic SB1 RV variations (see the work by Holl et al. 2023a), and are polluting the present sample (even if some post-filtering took place, see also Sect. 10.1.3). If, in addition, the close pair is rotating on the plane of the sky as is the case for astrometric binaries, the measured RV could be corrupted (see Sect. 10.3). This problem will be corrected for DR4.
@@
Besides, it should be noted that all decisions are taken in a statistical way; this means that the output is statistically correct but that some incorrect individual fits can survive near the borders where the decisions are taken (e.g. between an orbital fit and a trend solution for example).
@@
We recommend to the user of the catalogue to keep in mind the weaknesses and caveats described here above. We also strongly recommend, when possible, the use of the various quality factors accompanying each solution. If the catalogue is expected to be well-behaved from the statistical point of view, the situation could be more intricate for individual objects or particular set of objects. It remains that biases in the catalogue could be present from the point of view of stellar population studies. Such an analysis is beyond the scope of the present paper. For the time being, the genuine selection function of the catalogue is poorly known and any user should be aware of this and behave accordingly. The user is strongly encouraged to adapt the selection of solutions to their particular objective instead of using it blindly. They should combine the various delivered quality factors according to their aim. We advise a cautious approach. An example of interesting complex combinations is detailed in Bashi et al. (2022). To some extent, other information can be considered as quality factors (such as flags but, mainly, the number of points in the time-series ). They could also be included in the selection process.
@@
The global weak point of the algorithms could be the determination of the period even if here above it is finally shown that the success rate is at the very least 80% and is more often located in the range 85-95%. There is evidently at a very low number of transits, a difficulty to correctly determine the period that forced us to neglect data with less than 10. However, at slightly larger , the situation is not so simple. From the simulations described in Sect. 9.1, we conclude that a value of smaller than 17 is somewhat problematic (although it depends on the signal-to-noise ratio) for the period recovery. We also showed from the validation (see Sect. 9.2), that an overfitting phenomenon characterises part of the solutions associated to an in the range 10-15. One should further add that, at small , the periodogram has a larger propensity to produce isolated aliasing peaks (generating ambiguity due to aliasing). While interpreting the simulations, it appears that it is unclear if the determining parameter is the or the significance. According to these simulations, significance lower than 10-15 can be associated with difficulties to have the proper recovery rate. This seems to be confirmed by the validation in Sect. 9.4, in particular in the corner plot of Fig. 46. It is also clear, from Fig. 30, that the circularisation phenomenon is not outstandingly visible for significance much lower than 30. Therefore, we consider that depending on the studies performed, the significance should be cut between 15 and 25, or more conservatively at 30. If 10 is still affordable in some cases, the solutions characterised by a significance between 5 (our cut-off) and 10 are most probably hazardous to accept.
@@
The spectroscopic solutions could be merged with photometric or astrometric ones, and in these cases the resulting output is certainly very secure for the period. If the user of the catalogue only considers the sole results of the spectroscopy fits, it is advisable to restrict the sample to some kind of gold sample. This can be done by performing additional selections in the catalogue: for example, a larger number of transits, a higher significance, periods in a dedicated range (between ten days and , or allowing for going down to smaller periods but at the expense of a selection on the basis of a larger significance), period with small uncertainties, solutions with negligible phase gaps (although the gap values are not provided in the DR3 catalogue), etc…
In Appendix J, we present indicative but simple examples of extraction of a gold sample from the catalogue.
@@ 12 Conclusions
In the framework of the Gaia Data Release 3, we present here the first version of the Gaia catalogue of spectroscopic orbital solutions (SB-subcatalogue). We present 181327 objects with a solution of class SB1 and 202 objects with a solution of class SB1C. We also present objects exhibiting trends of variations of their RVs with time.
The present catalogue constitutes by far the largest collection of objects with a spectroscopic orbital solution, as can be evaluated by comparison with the notorious SB9 compilation171717https://sb9.astro.ulb.ac.be/. The improvement in number of concerned objects is more than one (almost two) order of magnitude. The present SB-subcatalogue has also the major advantage of being homogeneous from the point of view of both data acquisition (properties) and data processing, surpassing the characteristics of existing compilation. The only exception is the typical number of measurements that constitutes the individual time-series, which remains limited (at least for the present DR3 catalogue). The situation will undoubtedly improve for future data releases.
After a description of the data entering the NSS-SB1 processing, we describe in details the logic and structure of the related code. The general philosophy behind it is explicited along with the mathematical formalism used. We also describe, after the main processing, the filtering of the results that we performed to discard bad solutions and improve the catalogue. Examples illustrating the quality of the output are given as well as a few examples of detected problematic cases. A full validation of the output is performed both on the basis of internal consistency and on the basis of external existing catalogues from the literature. The characteristics of the catalogue are further analysed. The way the SB-subcatalogue is included in the Gaia Data Release 3 is also described. A section is included that scrutinises the weaknesses and provides caveats and recommendations for the scientific use of the catalogue.
@@
The epoch RVs constituting the time-series analysed here have not been published in the context of the Gaia Data Release 3. The orbital solutions and trends we derived are available since June 13, 2022, in the ESA Gaia DR3 archive accessible at https://gea.esac.esa.int/archive. The same results can also be found at CDS Strasbourg via the link https://cdsarc.cds.unistra.fr/viz-bin/Cat?I/357.

### v2

@@ 10.1.4 Further remarks
Although the post-filtering has been applied with great caution, it also happens that some problems persist. One of these cases is the following. Some objects exhibit a good SB1 or SB1C solution with a well established significant orbital semi-amplitude for Gaia. However, these solutions could sometimes be doubtful because the same object observed in the framework of the APOGEE DR17 survey (Abdurro’uf et al. 2022) or LAMOST survey (Cui et al. 2012; Bashi et al. 2022) appears to be constant in RV as detailed in Sect. 8.2.4. This anomaly has not been fully explained but it could be related in some cases to the presence of a nearby neighbour (see below). These objects have not been rejected. We encourage the catalogue user to be very cautious when using the present SB subcatalogue despite the fact that persisting problems should only represent a minority. All the remaining and identified problems should be corrected in the future DR4.
@@ 10.2 An explanation for Gaia DR3 3376949338201650112
In Sect. 8.2.4, we report the existence of SB1 solutions presenting a marked periodic variation, whereas ground-based data conclude to the constancy of the related object. Here below, we propose an explanation for this discrepancy by investigating in details the case of Gaia DR3 3376949338201650112. This example has the advantage of being enlightening and will allow us to correct the pipeline for future releases.
@@
The astrometric solution for this object is characterised by a ruwe of 25.35, pinpointing to a problem with the fit. Referring to the IPD parameters (image parameter determination; see e.g. Holl et al. 2023a), we remark that ipd_gof_harmonic_amplitude is equal to 0.164 which is certainly different from zero. This is a clue that the object is badly treated as a single object. The value of ipd_frac_multi_peak turns out to be 84 suggesting that the object is expected to be double or multiple on the plane of the sky: most probably a resolved or, more precisely, a partially resolved pair (according to the definition of Holl et al. 2023a). As explained by these authors, such pairs (or higher order) have a predisposition to suffer from scan-angle effects that induce a bias on the measured RVs that is function of the relative orientation of the scanning direction on the sky at the time of the observations. This is confirmed for Gaia DR3 3376949338201650112 by Fig. 50 where it becomes clear that the measured RVs are function of the scan angle. The parameter ipd_gof_harmonic_phase adopts a value of 103.4 ° that suggests that the line joining the two stars of the pair has a position angle on the sky of 13.4 °, after the application of a 90 ° correction as explained in Sect. 3.1.2 of Holl et al. (2023a). The semi-amplitude of the fake orbital solution is about 16.6 km s-1 which translates according to Eq. 2 of Holl et al. (2023a) into a separation of the pair of 114 mas. The existence of such a pair is confirmed by a preliminary run of AGIS (AGIS 4.1) linked to the future DR4 that actually detects two objects Gaia DR4 3376949342501945344 and Gaia DR4 3376949342501945472 that are separated by some 200 mas and are exhibiting an orientation of 9 °, thereby supporting our interpretation.
@@
In principle, the bias due to the scan-angle effect is intended to produce variability with a period either of 62.97 d or of 1 y. This property has been used in Sect. 10.1.3 to apply post-filtering. However, in the present case, the period is 0.657582 d which does not appear to be compatible and could shed some doubt on our interpretation. The spectral window associated to the DR3 RV time series corresponding to this object is given in Fig. 51. Besides the main peak at d-1 as usual, it exhibits an important peak at d-1 that is exceeding a value of 0.8; this is an indication that isolated aliasing could be important for this particular time series. An aliasing with d-1 explains the ambiguity between a period of 0.657582 d and that of 1 year. The combination of the aliasing phenomenon and of the scan-angle effect explains the existence of such fake SB1 objects in our SB subcatalogue: such objects were not post-filtered, since the post-filtering was based on the 62.97 d or 1 y timescales. We conclude that it would be more appropriate to include a statistical test inspired by graphs like the one shown in Fig. 50 rather than on a criterion based on the period. However, a true periodicity with these timescales will imply a similar behaviour of the RV dependency on the scan angle even if no scan-angle effect is present.
@@ 10.3 The case of astrometric binaries
In Sects. 10.1.3 and 10.2, we discussed the effect of astrometric close pairs on the determination of RVs. A close pair fixed on the sky could often generate a scan-angle effect and lead to corrupted RV values but also could generate fake orbital motions. An additional degree of complexity is reached if we consider astrometric binaries. Due to their binarity, these close pairs are not fixed on the sky and the main axis of the pair is rotating due to the orbital motion. Therefore, there is interaction between the above-mentioned scan-angle effect and the astrometric orbital motion itself. The situation still increases in complexity if the astrometric orbital motion and the spectroscopic RV changes correspond to two aspects of the very same orbital motion and thus have the same period. In this latter case, the correction presents a timescale equal to the phenomenon we are studying. The correcting scheme is explained in Sect. 3.3.1 of Holl et al. (2023a) and is also illustrated there with the (somewhat extreme) example of Gaia DR3 6631710607341412096. For details, we refer to this paper.
@@ 11.1 Description of the output
The present first catalogue of pure spectroscopic solutions is part of the Gaia Data Release 3. The orbital solutions (in class SB1 or SB1C as indicated in the entry nss_solution_id) are part of an auxiliary table named gaiadr3.nss_two_body_orbit. We output in this table the classical Campbell orbital parameters introduced in Sect. 3.1 for the SB1 class (, , , , , ) along with their estimated uncertainties (1). The same is done for circular orbit parameters (, , , ) introduced in Sect. 3.2 and corresponding to the SB1C class. Beyond the orbital parameters themselves, this table also contains some indicators of the quality of the solution (see Sect. 6.4). The flags associated to the solution (see Sect. 6.5) are also provided. The boolean mask bit_index defines the parameters that are actually fitted, and corresponds for the SB1 class (respectively SB1C) to the value 127 (respectively 31). In addition, is also provided the upper triangle of the correlation matrix of the parameters. It is expressed as corr_vec that contains the cross-coefficients folded into a vector and corresponding to the ordering of the adopted list of parameters. The first term of the vector gives the element (0,1) of the matrix, the element gives the element (0,2) of the matrix and element the (1,2) one. Elements , , correspond respectively to the matrix elements (0,3), (1,3), (2,3), and so on. This is thus a column-major storage scheme. The details of the output are described in Table 15 for the class SB1 and in Table 16 for the SB1C one. The table contains 181327 objects (with solutions) in the SB1 class and 202 in the SB1C one. The marked difference in the counting is due to our decision to consider eccentric solutions down to as small eccentricities as possible before going circular. This was described in Sect. 6.3.
@@
11.2 Weaknesses, caveats, and recommendations
@@
Another tricky case is that of true SB2 stars, where the two lines never fully deblend. They are appearing most probably in the SB1 channel. The RV variations are then due to the corresponding change of the full profile made by the blended lines. This has little effect on the period determination but the impact on the eccentricity and mainly on the semi-amplitude might be huge (Sect. 8.2.3). The related solutions are characterised by a too small value of and of the mass function. They could be considered as suspicious but we cannot have any certitude that all of these types of solutions are necessarily wrong.
@@
Instrumental effects are also present. The scan-angle effect notably induces periodic RV variations for constant stars that do not fulfil the assumptions used in astrometry (objects are assumed to be isolated and thus single). Close astrometric pairs could thus mimic SB1 RV variations (see the work by Holl et al. 2023a), and are polluting the present sample (even if some post-filtering took place; see also Sect. 10.1.3). If, in addition, the close pair is rotating on the plane of the sky as is the case for astrometric binaries, the measured RV could be corrupted (see Sect. 10.3). This problem will be corrected for DR4.
@@
In addition, it should be noted that all decisions are taken in a statistical way; this means that the output is statistically correct but that some incorrect individual fits can survive near the borders where the decisions are taken (e.g. between an orbital fit and a trend solution for example).
@@
We recommend to the user of the catalogue to keep in mind the weaknesses and caveats described here above. We also strongly recommend, when possible, the use of the various quality factors accompanying each solution. If the catalogue is expected to be well-behaved from the statistical point of view, the situation could be more intricate for individual objects or particular set of objects. It remains that biases in the catalogue could be present from the point of view of stellar population studies. Such an analysis is beyond the scope of the present paper. For the time being, the genuine selection function of the catalogue is poorly known and any user should be aware of this and behave accordingly. The user is strongly encouraged to adapt the selection of solutions to their particular objective instead of using it blindly. They should combine the various delivered quality factors according to their aim. We advise a cautious approach. An example of interesting complex combinations is detailed in Bashi et al. (2022). To some extent, other information can be considered as quality factors (such as flags but, mainly, the number of points in the time series ), which could also be included in the selection process.
@@
The global weak point of the algorithms could be the determination of the period even if here above it is shown that the success rate is at the very least 80% and is more often located in the range 85-95%. There is evidently at a very low number of transits, a difficulty to correctly determine the period that forced us to neglect data with less than 10. However, at slightly larger , the situation is not so simple. From the simulations described in Sect. 9.1, we conclude that a value of smaller than 17 is somewhat problematic (although it depends on the signal-to-noise ratio) for the period recovery. We also show, from the validation (see Sect. 9.2), that an overfitting phenomenon characterises part of the solutions associated to an in the range 10-15. We should further add that, at small , the periodogram has a larger propensity to produce isolated aliasing peaks (generating ambiguity due to aliasing). While interpreting the simulations, it appears that it is unclear if the determining parameter is the or the significance. According to these simulations, significance lower than 10-15 can be associated with difficulties to have the proper recovery rate. This seems to be confirmed by the validation in Sect. 9.4, in particular in the corner plot of Fig. 46. It is also clear, from Fig. 30, that the circularisation phenomenon is not outstandingly visible for significance much lower than 30. Therefore, we consider that depending on the studies performed, the significance should be cut between 15 and 25, or more conservatively at 30. If 10 is still affordable in some cases, the solutions characterised by a significance between 5 (our cut-off) and 10 are most probably hazardous to accept.
@@
The spectroscopic solutions could be merged with photometric or astrometric ones, and in these cases the resulting output is certainly very secure for the period. If the user of the catalogue only considers the sole results of the spectroscopy fits, it is advisable to restrict the sample to some kind of gold sample. This can be done by performing additional selections in the catalogue: for example, a larger number of transits, a higher significance, periods in a dedicated range (between ten days and , or allowing it to go down to smaller periods but at the expense of a selection on the basis of a larger significance), periods with small uncertainties, solutions with negligible phase gaps (although the gap values are not provided in the DR3 catalogue), and so on. In Appendix J, we present indicative but simple examples of extraction of a gold sample from the catalogue.
@@ 12 Conclusions
Within the framework of the Gaia Data Release 3, we present the first version of the Gaia catalogue of spectroscopic orbital solutions (SB subcatalogue). We present 181 327 objects that have been found to be characterised by class SB1 and 202 objects by class SB1C. We also present objects exhibiting trends of variations of their RVs over time.
The present catalogue constitutes the largest collection of objects with a spectroscopic orbital solution by far, based on a comparison with the notorious SB9 compilation181818https://sb9.astro.ulb.ac.be/. The improvement in terms of number of objects is more than one order (nearly two) of magnitude. The present SB subcatalogue has also the major advantage of being homogeneous from the point of view of both data acquisition (properties) and data processing, surpassing the characteristics of existing compilation. The only exception is the typical number of measurements that constitutes the individual time series, which remains limited (at least for the present DR3 catalogue). The situation will undoubtedly improve for future data releases.
After a description of the data entering the NSS-SB1 processing, we describe in detail the logic and structure of the related code. We reveal the general philosophy behind, along with the mathematical formalism used. We also describe, following the main processing details, the filtering of the results performed to discard bad solutions and improve the catalogue. We give a number of examples illustrating the quality of the output, as well as a few examples of detected problematic cases. A full validation of the output is performed both on the basis of internal consistency and on the basis of external existing catalogues from the literature. We further analyse the characteristics of the catalogue and also describe the way the SB subcatalogue has been included in the Gaia Data Release 3. We additionally scrutinises the weaknesses of the approach and provide the necessary caveats and recommendations for scientific applications of the catalogue.
@@
The epoch RVs constituting the time series analysed here have not been published in the context of the Gaia Data Release 3. The orbital solutions and trends we derived are available since 13 June 2022, in the ESA Gaia DR3 archive accessible at https://gea.esac.esa.int/archive. The same results can also be found at CDS Strasbourg via the link https://cdsarc.cds.unistra.fr/viz-bin/Cat?I/357. Appendices D to J are only available on Zenodo (see https://zenodo.org/records/13990211).